OpenStudy (anonymous):

This is for "manupulation of formulas" a*b/c = d*e/f ...I need to solve for the following: solve for d solve for e solve for f solve for c

OpenStudy (anonymous):

This should really be in mathematics but here goes anyway: d = fab/ce e = fab/cd f = dec/ab c = fab/de